R/V ENDEAVOR Cruise 302

to Georges Bank

9 - 22 June 1997


We gratefully acknowledge the assistance of the Captain, Officers, and crew of the R/V Endeavor and of Jan Szelag, Marine Technician. Their assistance contributed greatly to the success of this cruise. Thanks also to C. Flagg and M. Dunn at Brookhaven National Laboratory, who assisted in preparing the ADCP data for this cruise report, and J. Manning (NMFS) and J. Churchill (WHOI) for their assistance with the drifters.

This report was prepared by Carin Ashjian and Cabell Davis. This cruise was supported by the National Science Foundation.

1. Introduction

The goal of the U.S. GLOBEC program is to describe how biological and physical characteristics interact to control the abundance and production of marine populations, including marine fish. It is recognized that important controls of populations abundances operate during the early, usually planktonic, life stages of the organisms.

The U.S. GLOBEC Georges Bank program focuses on the biological and physical mechanisms regulating the abundance of cod and haddock larvae and the dominant zooplankton species, Calanus finmarchicus and Pseudocalanus spp., upon which the larval fish feed. The program combines field sampling, data analysis, and modeling.

The field sampling includes broadscale sampling (conducted monthly from January to June each year) to quantify population abundance and distribution over the bank and process studies conducted every other year which a) target specific physical mechanisms that affect population abundances on the bank and b) obtain traditional life history rate measurements. In 1995, the GLOBEC Georges Bank program focused on the development of vernal stratification and its effect on the growth and survival of the target species during the process cruises. In 1997, the program focused on the source, retention, and loss of plankton populations, with the dominant physical process being the transport of plankton on and off the bank by ocean currents.

1.1 Purpose of Cruise

As part of this effort, the purpose of the present cruise was to use the Video Plankton Recorder (VPR) together with the ship's hull-mounted Acoustic Doppler Current Profiler (ADCP) to quantify the flux of the copepods Calanus finmarchicus and Pseudocalanus spp. through the "retention/loss" region in the southwestern corner of the Bank. Previous work has suggested that this region is a critical location for either the loss of plankton populations to the southwest in alongshelf currents or retention of plankton populations in currents that turn to the north and recirculate around the Bank (e.g., Davis, 1987). Hence, this southwest corner of the bank is likely to be the major region where the copepods (and other plankton) are either lost or retained on the bank. The presence of recirculation in this region is thought to be related to the onset of seasonal stratification which begins typically in April. ARGOS drifters from previous GLOBEC years were found to be more likely to recirculate around the bank after the onset of stratification (Limeburner et al.). The present cruse was the second of two retention/loss VPR cruises conducted during 1997; the first cruise was conducted using the R/V Cape Hatteras (CH0697) in April during the onset of stratification.

The flux of copepods through this region may be determined from their concentration and the current magnitude and direction. In this application, the VPR serves as tool for measuring the concentrations of plankton and the magnitude of other variables (temperature, conductivity, fluorescence, beam attenuation, and incident light) with high along-path resolution (cm to 100s of kilometers). The sampling rate presently is 60 Hz for video and 3-6 Hz for the ancillary parameters. The product of these copepod concentrations and the advection rate in this region yields an estimate of the flux of these animals.

1.2 Cruise Overview (See Appendix 1)

During the cruise, we conducted several types of sampling to measure the copepod concentrations and water flow through the retention/loss region: 1) 24-hour towyo sections along both the N/S and E/W Schlitz mooring lines to measure the tidally resolved flux into the region at the inflow end and out of the region as recirculation, 2) a 24-h diel vertical migration study which consisted of continuos surface - bottom VPR towyos conducted along a track which ran around 10 and 30 m drogued GPS/ARGOS drifters (two drifters were launched at each of the 65, 75, and 85 m isobaths), and 3) a zig-zag "grid" E-W, N-S covering the retention loss area bounded to the east by the N/S Schlitz mooring line and to the north by the E/W Schlitz mooring line while tracking the paths of the 10 and 30 m drogued drifter pairs that had been deployed at the 65, 75, and 85 m isobath along the N-S Schlitz mooring line (Figure 1; Table 1). Following completion of these objectives, we moved to the north of the E-W Schlitz mooring line and towyoed along a zig-zag path across and along the northern edge of Georges Bank to document influx of plankton populations onto the Bank from both the Great South Channel and in water recirculated from the south. The calm weather afforded us the opportunity to conduct a survey of the Great South Channel region to document copepod abundances and patchiness in that region; we followed a chevron shaped cruise track along E-W towyo transects between the shipping lanes to the west and the Bank to the east. Our final activity was to describe the shape and persistence and the diel vertical migration patterns of a patch of plankton located at approximately 41° 22' N, 68° 55' W by towyoing the VPR first in a star pattern and then around an ARGOS drifter.

The cruise was highly successful. We accomplished all of our objectives ahead of schedule and in addition were able to conduct the surveys of the northeastern corner of Georges Bank and the Great South Channel and the patch studies. No days were lost to weather or equipment failure.

2. Cruise Narrative (in local time except where noted)

June 9, 1997 - Monday

1415 The R/V Endeavor arrived in Woods Hole for loading of cruise gear. We loaded computer equipment and supplies during the afternoon. The power pack for the winch also was loaded onto 01 deck of ship. The winch and VPR were still at Dyer's Dock. Upon termination of the fiber optic cable, it was discovered that all three fibers were broken in at least one location. We arranged to have an instrument which detects breaks and approximate locations in fibers sent to WHOI to arrive at 10AM on June 10. Also, more work on the winch power pack was necessary to make it usable which was conducted while at the dock. The VHF and Gonio antennas were installed during late afternoon/evening for use in drifter tracking. Endeavor remained at the WHOI dock for the night.

June 10, 1997 - Tuesday

We continued setup on board ship. The drifters and drogues were loaded and secured on the main deck. The winch and fish arrived at ~1500 and were loaded together with the remainder of the VPR equipment from Dyers dock. The winch was welded in place by ~1900 and the winch-power pack connections established. Tie down of interior equipment finished by evening. Scott Gallager and Andy Girard worked during the evening/night to terminate the fiber optic cable and to test the VPR.

June 11, 1997 - Wednesday

All hands aboard at 8AM for planned departure. We delayed departure for ~1.5 hours to complete setup of VPR and to conduct a dock "dunk" test of fish.

We departed WHOI at 1345 (UTC). Our first science activity was to test the VPR in Vineyard Sound in order to check the pitch and roll of the fish during towyoing and at high (10kts) and low speeds. VPR 1 (test tow) was conducted in the Sound at 1413 (UTC) (Table 3). The fish rolled a bit during the first deployment. We brought the fish on board and exchanged the port trim tab for a slightly longer piece; the fish flew nicely following the adjustment.

Departed Vineyard Sound at 1601, heading to first station on Georges Bank. Estimated arrival time is June 12, 0000. Drills at 1300.

June 12, 1997 - Thursday

We arrived at first station/waypoint at 0055 local. The VPR was deployed for first tow (VPR 2) of the science plan at 0539 (UTC) (Table 3). Our first activity was a 25 hour survey between waypoints 1 and 3 (Table 1), towyoing continuously along the transect line. We planned to do 12 transects total. The ship speed was ~ 9 knots. Leg 1 was complete at 0817 (UTC). We had to deviate from the transect at 1005 (UTC) in order to avoid 2 scallop boats with gear deployed. Leg 2 was complete at 1021 (UTC).. Leg 3 was complete at 1218 (UTC). We observed high concentrations of Chaetoceros socialis colonies and Obelia medusae along the transect, especially at the northern edge, in both deep and shallow water. At southern end of transect, the water column contains mixed layer, the cold pool, and warm salty water under the cold pool. C. socialis was found in the mixed layer and down to the cold pool/warm water interface. Not much was seen in the warm salty water at depth. At the northern end of the transect, there was a well mixed water column with Chaetoceros throughout and high concentrations of medusae and teardrop shaped marine snow near the bottom. Leg 4 was complete at 1436 (UTC), Leg 5 at 1702 (UTC), and Leg 6 at 1852 (UTC).

1953 (UTC). The VPR was brought on board (end of VPR 2) at the end of Leg 6 of the 24 hour tidal section to adjust the trim as the roll was getting rather high in the negative direction. Approximately 2 cm was cut off the bottom of the starboard trim tab to bring the roll more positive. During adjustment of trim tabs, Andy Girard noted that the lens cover for Camera 4 was foggy. Closer examination revealed that there was about 2 ml of water in the camera housing and that the glass window was cracked. It was hypothesized that thermal contraction of the aluminum cap around the window caused the window to crack. The window and cap were replaced with an older window/cap housing. The camera lens was dried in the toaster oven at 100 °F. The VPR was deployed for VPR 3 at 2145 (UTC) at the location where VPR 2 had terminated (beginning of Leg 7) (Table 3). Leg 7 was complete at 2253 (UTC).

June 13 , 1997 - Friday

We continued along N-S “tidal section”. Leg 8 was complete at 0115 (UTC). Obelia, chain diatoms, and Chaetoceros were seen in upper water column above the cold pool. Copepods were seen below the Chaetoceros layer and teardrop shaped marine snow at depth. Leg 10 was complete at 0317 (UTC), Leg 11 at 0817, and Leg 12 at 0950 (UTC). The VPR was brought aboard at the terminus of Leg 12, ending VPR 3 (Table 3).

Following the end of VPR 3, we moved to a transect line 2 miles to the west of the sampled transect line. This new line was located approximately 1 mile west of the Schlitz mooring line and ran parallel to our sampled line. We planned to deploy two drifters (drogued to 10 and 30 m) at each of three isobaths (65, 75, and 85 m). We moved down the line until we reached the 65 m isobath. During transit, and after arriving at the site, we checked out the drifters and strobes to ensure that they were functioning properly. We selected two of Ted Durbin's drifters for this site as they do not have VHF communication and might be detected at a shorter range than those with VHF. The drifters checked out fine; all three were detected using the Gonio box. The strobes, however, were in sorry shape. Scott Gallager and Andy Girard worked heroically but were able to revive only 5 strobes in total. Two drifters (#7201, drogued to 10 m, and #7234, drogued to 30 m) were deployed at the 65 m isobath at 1206 and 1209 (UTC), respectively (Table 3). After insuring that the drifters had maintained communication following immersion in water, we proceeded along the transect line to the 75 m isobath where we deployed 2 more drifters (one of Ted Durbin's and one of Jim Manning's) at 1313 (UTC) and 1314 (UTC) (Table 3). The drifters were drogued to 10 m (#34) and 30 m (#7200). Following deployment and check out of the signals, we moved further south to the 85 m isobath where we deployed two VHF equipped drifters (Manning), at 1348 (#36; 10 m) and 1349 (UTC) (#21; 30 m) (Table 3). Of these two drifters, one (#36; 10 m) was drogued using a small, ~18" diameter drogue. We had ordered additional drogues, however we did not receive them prior to sailing so resorted to using a small drogue borrowed from R. Geyer. The second drifter (#21, 30 m) was not equipped with a strobe.

VPR 4 commenced following drifter deployment at 1406 (UTC) ((Table 3). The purpose of VPR 4 was a ~24 hour survey around the 3 drifter pairs to examine diel changes in the vertical distribution of taxa in these water parcels. Drifter tracking was accomplished through a combination of Gonio tracking and VHF tracking using data inputs into Matlab via the Winwedge program. Matlab routines for this purpose originally were developed by Craig Lewis and were modified by Jim Manning. The routines interpreted the Gonio string, displaying the latest positions of the drifters #7200, 7201, and 7234. Unfortunately, a seventh drifter on the ship was not transmitting meaningful Gonio data and could not be used in calculating range and bearing of each drifter from the ship. A similar routine interpreted the VHF data strings for the 3 drifters equipped with VHF (#21, 34, and 36). For this method, the drifter on board the ship was functioning properly and the locations of the drifters relative to the ship could be calculated. However, we are not receiving data from the ARGOS center for the three Durbin drifters. We contacted R. Campbell (URI) and J. Churchill (WHOI) for assistance.

June 14, 1997 - Saturday

We continued VPR 4 and tracking drifters. It was foggy but calm . All 6 drifters were giving good fixes. Torrential rains and lightening occurred during 0200-0400 (UTC). Drifter #36 (southern deployment, 10 m drogue, Rocky Geyer's small drogue) had either experienced a huge N-S excursion in its tidal ellipse or had lost its drogue; the drifter was headed straight for Waypoint #1 and out of the grid (Table 1).

1139 (UTC) The C. socialis patch at 12-15 m seemed to remain at constant depth during both day and night. The fluorescence above (0-12m) varied throughout the diel cycle; fluorescence was low during the day but increased to higher levels in the early evening and early night (cell doubling?). After midnight, fluorescence in the upper 12 m appeared to decline. Could this be because of grazing during the night? Also noted decomposing / fragmented C. socialis colonies between 35 and 50 m.

1150 (UTC) Jim Churchill called. He had contacted ARGOS and they were receiving signals from the Durbin drifters. Apparently the subscription to transmit for these three drifters had expired. Jackie Mason at ARGOS will fax the positions to us a couple of times a day over the weekend and also to Jim. Jim will e-mail them to us as well. Next week Jim will arrange to have the drifter positions e-mailed to us on the ship regularly.

1306 (UTC) Drifter #36 continued to go south and was leaving the study area!

1421 (UTC) End of VPR 4; we brought the VPR on board at the completion of the 24 hour diel study around the drifter (Table 3) and headed towards the latest position of Drifter #36.

Scott checked out the VPR while on deck. The cable was kinked between strain relief and fairing and there were marks on the fin where it obviously hit the block. No strands were broken at this time. The secondary armor did not appear to be damaged other than the kink (2 ° armor is visible through spreading of threads in 1° armor). Scott retaped the missing piece of fairing, approximately 6 ft. were lost from the area around the kink. The strain relief was marred from the block but not cut through. There were tiny drops of water on the inside of the camera window. The outside was cleaned but not the inside. Both the wire and the camera window should be checked after about 6-8 hours more of fast (8kt.) towing.

1532 (UTC) Recovered Drifter #36 (Table 3). The drogue was intact; the drifter was just heading southwest out of the study area at a rapid rate.

1556 (UTC) We started VPR 5 at Waypoint #1 of the zig-zag survey (Tables 1 and 2; Appendix 1). The towing speed was 8 kts.

2055 (UTC) Four of the drifters had moved generally southwest from the original deployment location. The fifth (#7234, drogued to 30 m and deployed at 65 m isobath) had moved basically to the west.

2226 (UTC) We terminated VPR 5 at the end of Leg 2 (Tables 1 and 2; Appendix 1) and brought the fish on board in order to examine the kink in the cable. The kink in the cable is the same as before (no progression or breakage). Camera 4 still had very small droplets of water on the inside of the window with no change in appearance of the droplets; we decided not to change or reset the "O" ring.

2239 (UTC) VPR # 6 began at Waypoint #12 (Table 3, Appendix 1), at the beginning of Leg 3. The weather is mostly sunny with low seas.

2338 (UTC) Received fixes (Gonio and VHF) from four out of five drifters. No fix was received from #234 as we were too far south from its last position.

2359 (UTC) We saw extremely high concentrations of Chaetoceros in the upper water column at the 70 m isobath as we were heading north on Leg 3.

June 15, 1997 - Sunday

0005 (UTC) Exceptional sunset, enjoyed by all. This bodes well for tomorrow's weather. We received fixes from all 5 of the drifters at 0020 (UTC) and again at 0040-0100 (UTC).

At 0213 (UTC) we noticed that the date on Balanion, the primary data logging computer, had not changed at midnight to be June 16. Therefore partway through both the zip and NFS log files 06142238.y97 the time changed from 2340 to 0000 but the date did not change. This will be corrected during post-cruise processing.

022350 (UTC) We started turn from Line 3 to Line 4 at Waypoint #5, holding VCR at 10 m. We closed the NFS and zip logging files to reset the date on Balanion, restarted with correct date.

0348-0350 (UTC) We are at the northern edge of Leg 4. Here the Chaetoceros are thick in the water from the bottom (66.5) to 12 m , with much lower concentrations in the upper 12 m. At 0450 (UTC) we experienced a whiteout on the video because of high Chaetoceros concentrations at < 30 m depth.

0547 (UTC) We started the turn towards Waypoint # 6 at beginning of Leg 5. At 0918 we turned south at Waypoint # 6 to Leg 6. We received drifter positions on Gonio at times ranging from 0635 (UTC) to 0715 (UTC).

1200 (UTC) The winch was overheating. The hydraulic fluid was much too hot (> 120 °F) and the winch control valve in the doghouse also was very hot. The situation was considered by several engineers, Andy, and Scott. Andy used cold seawater to cool down the winch motor which should cool down the fluid to appropriate temperatures as well.

1253 (UTC) This was a beautiful day, 1-2 ' swells, sunny and warm. We were heading towards the end of Leg 6 (Waypoint 10). At 1323 (UTC) we started the turn from Leg 6 onto Leg 7, towards Waypoint #7 to the north. The temperature on the hydraulic fluid tank was 015 °F. Andy's hose has worked! He reduced the volume of spray on the winch motor. At 1330 (UTC) the turn was complete and we resumed towyoing. As we headed north, there were thick bands of C. socialis in the water column associated with cold water and the Slope - Shelf front. The Chaetoceros were not abundant in the upper 10-12 m of the water column. At 1750 (UTC) we turned onto Leg 8. The ship slowed to 6 kts at 1902 (UTC) to permit a freighter to pass.

We noted convergent slicks with flotsam to be numerous in the region of 40 ° 50.63 N and 68 ° 43.96 W at 20:22 (UTC). At 2124 (UTC) we reached waypoint #4 and turned onto Leg 9, heading SSW. There wee teardrop shaped marine snow particles near the bottom and Pleurobrachia and medusa in the upper 10 m of the water column. We received fixes on 3 drifters at 2349 (UTC).

June 16, 1997 - Monday

0000(UTC) Once again, Balanion (data logging computer) did not execute the date change correctly. We halted the logging program to change the code in the program. Scott removed the "date jam" feature of the program. In its latest configuration, the program jams time and date to GPS only once at the beginning of a log file.

0056 (UTC) The water was thick with Chaetoceros from 25 m to the bottom at the 75 m isobath on the western end of Leg 9 of the zig-zag. At 0113 (UTC) we turned from line 9 onto Leg 10, at waypoint #8. We received good fixes from 4 drifters at ~0200(UTC) (#s 21, 34, 200, 201 and from all of the drifters at both ~0400 (UTC) and a ~0630 (UTC).

We turned onto Leg 11 at 0806 (UTC). Again, we got good fixes from four drifters. This continued success with the drifters was comforting as we did not have any information from ARGOS. So far, the VPR had been running for 126 hours this cruise. The hydraulic fluid in the winch continued to maintain an appropriate temperature, with the hose cooling the winch motor. We received good drifter fixes at 0935 (UTC). We turned onto Leg 12 at 1201 (UTC). Good fix on drifter #21 at 1415 (UTC). At 1551 (UTC) we completed turn onto Leg 13, the last leg of the zig zag survey. The remainder of the survey was uneventful, with completion of the survey and VPR 6 at 1944 (UTC) (Table 3).

We had spoken with Jim Churchill at 0730 and again at about 0930 regarding the failure to receive fixes from the ARGOS service on Jim Manning's e-mail. He had checked out the problem. Apparently Service ARGOS had some problem with their automatic distribution service (ADS) over the weekend and was unable to transmit information until Monday. The situation was to be rectified on Monday. We logged in to Jim Manning's e-mail on Monday afternoon (local) and several times on Monday evening (local) and obtained ARGOS service center fixes on all of the drifters.

We had launched six drifters originally. One of them (#36) had been recovered following the 24 hour tidal section because it was moving out of the study region to the SE. Four of the remaining drifters were tracking W-SW, grouped quite closely together, and it was evident at this time that the drifters would continue to move in this direction and exit the sampling region within 24 hours. We decided to recover these four drifters (#s 21, 34, 200, and 201). The sixth drifter, #734, was located to the NE of the group of four drifters. It was not clear whether this drifter would move to the north and towards the Great South Channel or track to the west as indication of loss from the Bank. We decided to leave this drifter in while we conducted the 24 hour tidal section across the northern end of the study grid. Since we were able to get information from the ARGOS service center regarding the position of the drifters, we were a bit more comfortable leaving this drifter in for a longer period.

We recovered four drifters (#s 21, 34, 200, 201) between 2202 and 2345 (UTC). All four were intact and appeared to be functioning properly. Data retrieval from two (#s 21, 36) was successful. The data recovery operation was terminated on account of rain and darkness.

June 17, 1997 - Tuesday

At 0005 (UTC) we were en route to the eastern end of the Northern E-W tidal section. We discussed our options for the remainder of the cruise and decided to conduct the tidal section, then recover the last drifter (#234), and then move up to do a survey of the western side of the Great South Channel to document the role of eddies and circulation in advection plankton populations on or off the Bank in that region. Following this survey, we hoped to conduct a survey of a copepod or pteropod patch in the GSC. We planned to mark the patch with one of the ARGOS drifters and then survey around it. We anticipated the arrival at the terminus of our tidal section at 0300 (UTC). We have been logging locations of the last drifter (#234) as we pass near its location.

The VPR was checked out during the period between the end of VPR 6 and the start of VPR 7. There was no deterioration of the wire in the region of the kink. The window on the Camera 4 tube still had some small drops of what appeared to be water or perhaps silicon from the 'O' ring.

VPR 7, a 24 hour E-W tidal section at 8 kts. across the northern region of the study area and just to the north of the Schlitz mooring line, started at 0311 (UTC) (Table 3). The weather was foggy but not cold, with a light wind. We completed Leg 1 at 0516 (UTC), Leg 3 at 0916 (UTC), and Leg 2 somewhere in between these two times. Leg 4 was completed at 1132 (UTC). During Leg 5, we started to make note of ADCP velocities. At 1249 (UTC) we were located at the 68 m isobath, at 40° 50.12, 68° 47.9. The ADCP currents were primarily to the north (85 cm/sec at surface, 75 cm/sec at bottom) with little eastward component. At 1257 (UTC) noted that saw Chaetoceros socialis from the bottom up to about 36 m, the Chaetoceros thinned a bit at 30 m but were present up 'til 10 m when the colonies virtually disappeared. Leg 5 was complete at 1333 (UTC). Held the VPR at 10 m during the turn. There were only a few Chaetoceros at this depth (10 m) and location (71 m isobath, Waypoint #6). The ADCP velocities had diminished at this time or location to 75 cm/sec N, 10 cm/sec E at the surface and 60 cm/sec N, 10 cm/sec E at the bottom. The bottom topography was quite irregular. Leg 6 started at 1339 (UTC). At 1348 (UTC) we noted that there was just marine snow and rod diatoms at 46 m, with no Chaetoceros. The ADCP velocities had dropped way down (20 cm/sec N, 10 cm/sec E). At 1356 (UTC) we noted that there were C. socialis in the bottom layers of the water column. By 1353 (UTC) the ADCP currents had reversed direction and were still low (10 cm/sec S, 25 cm/sec E). However, there were C. socialis found throughout the water column below 10 m. ADCP velocities had increased to 25 cm/sec S, 20-25 cm/sec E by 1405 (UTC). C. socialis was still present at 36 m depth. At 1410 (UTC), there were still C. socialis in the water column at 40 ° 49.3, 68° 49.9 W, even though ADCP velocities were to the south-east at 35 cm/sec S, 30 cm/sec E (surface) or 15 cm/sec E (bottom). There was a increase in AGC on the ADCP, which translates into a peak in backscatter intensity, near the bottom. It was interesting that there were still C. socialis in the water column...were these colonies that had been advected to the north with the northern flow which were being advected back to the south? At 1447 (UTC) the bridge noted that the ship was being set to the SE pretty well. By 1505 (UTC) there were no C. socialis at 44m. ADCP velocities were firmly to the southeast, with 75 cm/sec S and 35 cm/sec E. The presence of C. socialis in the water column during flow to the SE suggested that we were seeing the return to the south of a blob or parcel of water, perhaps the edge of a front, which contained C. socialis and had been advected to the N with the northward flow. After the reversal of the flow to the SE, the C. socialis were still present at our location for some time as the "southern type" water moved back to the south. Subtle changes in temperature and salinity also were noted with the presence of C. socialis.

By 1511 (UTC) there were no C. socialis in the water column. Fluorescence was high at all depths, probably because of the high concentrations of diatom rods. During the periods when Chaetoceros was present, fluorescence had been reduced at depth relative to levels observed at this time and also much reduced in the upper 10 m where few Chaetoceros or rod diatoms were observed. The water containing Chaetoceros was associated with water that was warmer and fresher than that found in the "northern" type water.

Leg 6 was completed at 1532 (UTC). We held the VPR at 10 m during the turn onto Leg 7. ADCP velocities were to the SE (50 cm/sec S, 25 cm/sec E, ~160°, 55.9 cm/sec). Leg 7 began at 1538 (UTC). Slack water was at 1655 (UTC). Leg 9 began at 1951 (UTC) and Leg 10 at 2145 (UTC). at 2205 (UTC) we saw high abundances of Chaetoceros at the western end of Leg 10. However the colonies appeared to be decomposing and old. Oithona and Centropages were plentiful and hydroid medusae were abundant at < 40 m. Chaetoceros dropped off as we moved east. Turned to go back to the west on Leg 11 at 2345 (UTC).

June 18, 1997 - Wednesday

We began to see Chaetoceros again at 0045. Abundances increased until we were back in the Chaetoceros "patch" at 0135 (UTC). The real time plankton identification routines were working and we could see the Chaetoceros on the real time plot! The Chaetoceros was present in the colder (~8.5 °C) water below ~20m. The temperature of the water in the upper 20 m was ~11°C. Turned onto Leg 12 (last leg) at 0158 (UTC) and were heading to the east. C. socialis was still present in the water with ADCP velocities showing the current to be 15 cm/sec to the north in the upper water column. At 0248 (UTC) we started to pick up a SSE current, according to bridge observations. By 0300 (UTC) there were few Chaetoceros observed in the water column and the Chaetoceros had disappeared by 0318 (UTC). We reached the end of Leg 12 and retrieved the VPR at 0353 (UTC). The VPR was on deck at 0355 (UTC) (Table 3).

We headed south to retrieve the last deployed drifter (#234) at 0402 (UTC). Although there was heavy fog, we were able to locate and retrieve the drifter quickly at 0230 (UTC) following a 2 hour steam. We then headed back to the eastern end of the E-W tidal section to begin the zig-zag survey of the western side of the Great South Channel.

The purpose of the zig-zag survey was two-fold. First, we wanted to document the distribution of plankton in the GSC and the potential transport of plankton on or off the Bank in this region. Secondly, we wanted to identify the location of a plankton patch which we could map using the VPR following the zig-zag survey. The VPR was deployed at 0837 (UTC) for VPR 8 (zig-zag survey) and was towed at 8 kts (Table 3). In the initial legs of the survey (Legs 1, 2, and 3) we saw abundant marine snow and chain diatoms and few copepods. However, once we started moving out into the deeper water (> 100m) in the center of the GSC we started to see Calanus at depth. Also, the chain diatoms were absent from the water column. We saw whales along the end of Leg 3 and the beginning of Leg 4; the whales had fins on their backs and were leaping out of the water and were identified tentatively as humpbacks or fins. Leg 2 began at 1109 (UTC), Leg 3 at 1326 (UTC), and Leg 4 at ~1620 (UTC). Whales were sighted at 41° 17.77, 68° 55.27 at 1545 (UTC), 41° 19.6, 68° 58.2 at 1603 (UTC), and at 1630 (UTC) (no position recorded). We slowed the ship speed from 8 to 6 kts. at the turn onto Leg 4 at 1620 (UTC) in order to get better horizontal resolution of the water column (more towyos/nautical mile). Another whale was sighted at 1722 (UTC) at 41° 20.82, 68° 51.224 and a sunfish was observed alongside the ship at 1724 (UTC). Several shipwrecks were know to be located along our transect path so we reduced the depth to which the VPR was towed to 30 m off the bottom at 1834 (UTC). We resumed sending the VPR to 15 m off the bottom after passing the locations of the shipwrecks. Leg 4 was finished at 1940 (UTC). Along Leg 5, we slowed the ship speed to 6 knots in order to get the VPR to near bottom (120 m). The remainder of the tow was conducted at 6 kts in deep water and at 8 kts in shallow water (<120 m). A close encounter of the interesting kind with a fishing vessel occurred at 2241 (UTC); we had passed out of the area in the GSC closed to fishing. There was quite a congregation of fishing vessels just to the north of the boundary of the closed area (“just over the fence”). Saw Calanus at various depths along Leg 5. We began leg 6 at 2304 (UTC) and continued to see Calanus.

June 19, 1997 - Thursday

This day began with rain and lightning at 0029. The weather made it particularly uncomfortable to exchange personnel in the doghouse. Camera 4 on the VPR was fogging when the fish was in the top 10 m (warm water). The weather continued to be adverse, with squalls of heavy rain moving over the ship. We retrieved the VPR at 2241 (UTC), ending VPR 8 (Table 3).

The next activity was to conduct a chevron shaped survey of the GSC, with transect lines lying perpendicular to isobaths (as was done in SCOPEX) (Table 2). We hoped to identify copepod and pteropod patches with this survey and, if we identified such a patch, map it out in 3-dimensions using a sort of dynamic mapping technique. The chevron consisted of four transects which extended from Georges Bank on the east to just outside or just at the western boundary of the GSC closed fishing area. The western side, although the most probable location of cold fresh water advecting from the north and also of high abundances of Calanus and pteropods, was difficult logistically. This region was just outside the closed area (“over the fence”) and also was just to the east of the shipping lanes. Hence, high numbers of fishing vessels and also high speed shipping traffic was expected in this region. We had been beset by a heavy fog for days which made any kind of irregular surveying difficult in a high traffic area. Therefore, although we were likely to encounter high abundances of animals at the western side, the weather and traffic conditions combined were too dangerous to permit a patch survey at that location and time.

Scott Gallager worked to rectify the fogging of Camera 4 during the transit from the end of VPR 8 to the start point of VPR 9. Despite heating the lens in the toaster oven, fogging was still a problem. Furthermore, Camera 4 had to be re-aimed several times and the VPR was placed in and out of the water several times at approximately 0647 (UTC). VPR 9 finally commenced at 0659 (UTC). The first transect (most northern) was conducted from east to west across the GSC. Calanus was observed at various depths across the first transect but dense Calanus were not observed until just before the end of the transect, near the western side of the closed area, at 1126 (UTC). High abundances of Calanus were observed throughout the turn to the south in the region “over the fence” and to the east of the shipping lanes. The VPR was held at 10 m throughout the transit from the end of Line 1 to the beginning of Line 2. We turned onto Leg 2 at 1236 (UTC) and headed east at 8 knots, conducting VPR towyos between 5 and 60 m. Calanus was observed in the upper water column but dropped off in abundance by about 40 m. Few were seen in the upper 10 m.

At 1253 (UTC) a very strange event occurred; during haul back from 60 m the VPR suddenly “rocketed” to the surface. Both pitch and roll were pegged at 45°. It was quickly ascertained that the fish had flipped over onto its back and was surfing on the fin. There was no apparent reason for the flipping out of the VPR. We retrieved the VPR to examine it and determine if any damage had occurred. The fish was intact and was re-deployed immediately without terminating the logging programs.

We were still in heavy fog with the fog horn sounding at regular intervals at 1259 (UTC). We were towyoing the VPR between 5 and 60 m at 8 kts. As we moved east across Leg 2 we moved out of the high abundances of organisms that had been seen along the western edge of the survey region. Calanus were found in the thermocline between the cold salty water at depth and the warm fresh water at the surface. The peak in abundance of Calanus was coincident with both high fluorescence and high attenuation. Abundances of Calanus decreased along the transect towards the east until the central point of the chevron where the abundances increased again. Unfortunately, Xiaoou’s real-time copepod identification program had fallen behind temporarily so we were unable to utilize it to track the patch. At the central point we conducted at deep towyo to 15 m off the bottom, slowing the ship speed to 6 knots. We kept the ship speed at 6 kts. thereafter as it appeared that we were in a copepod patch. Calanus was located at the interface between the cold (<10 °C) salty water at depth and the warm (11-13 °C) fresh water in the upper 10 m. We also started seeing pteropods.

We appeared to move out of the copepod patch at 1449 (UTC), although we did continue to see Calanus but in lower abundance. and also pteropods. We picked up the speed of the ship to 8 knots at 1415 (UTC). As we moved to the eastern end of Leg 2 we started seeing chain diatoms. By 1537 (UTC) there was a veritable snowstorm of chain diatoms and also quite a few medusae. We turned onto the transit between Legs 2 and 3 at 1557 (UTC) and held the VPR at 10 m for most of the transit. At 1644 (UTC) we turned onto Leg 3, moving to the west at 8 knots and towyoing the VPR. At the turn onto Leg 3, saw a Coast Guard cutter on the radar, lurking nearby.

We slowed the ship to 6 kts at 1713 (UTC) in order to do some deep VPR tows across the center of the chevron survey. We were seeing numerous copepods in the upper portion of the water column, however the Calanus appeared to be located below the upper, warm layer in association with the thermocline and with the fluorescence maximum. At 1916 (UTC) the VPR had a close encounter with the block, which terminated operations for repairs. We brought the VPR on the deck at 1927 (UTC), ending VPR 9 (Table 3). The stabilizer fin was bent to the side with bent supports and a broken tail bracket that supports the light sensor. We were able to bend the stabilizer fin back into its proper orientation and remount the light sensor. During the time on deck we also removed and cleaned/dried Camera 2. The window of the Camera 2 housing cracked upon removal (too tight?) and so the entire aluminum cap was replaced with a spare cap. Also removed Camera 4 and reset the f-stop and alignment; we were very pleased by the image.

We redeployed the VPR at 2140 (UTC) as VPR 10 (Table 3). The purpose of VPR 10 was to locate a suitable patch or high concentration of animals for a high resolution study. wee headed to the NE from the south of Line 3 of VPR 9 to a point NE of the apex of Line 2 where we had noted high concentrations of Calanus during VPR 9. We had been seeing pteropod egg sacs but few pteropods. Now we were seeing high concentrations of pteropods at depth (60-78 m) at 2304 (UTC). The roll had gone positive, which is advantageous as the fish rolls away from the ship wake when near the surface.

We arrived at the end of our transect line at around 2345 (UTC) when we noted a Calanus hot spot at 41° 23.5, 68° 57.5. We started surveying the region using a “star” pattern where the ship travels for 5 nm in one direction, then turns 135° to the right of that course and travels for another 5 nm and then continues the pattern of turning and traveling 5 nm.

June 20, 1997 - Friday

The evening was foggy but calm as we embarked on the hunt for the elusive Calanus mother lode. We continued to follow the star pattern, towing the VPR from 0-30 m. We saw considerable variation in abundance of copepods both with depth and with horizontal distance. We documented the relative abundances of the animals using the roi identification routines on the Sun and SGI workstations. Xiaoou Tang’s routines were working nicely. It was helpful and exciting to see the data plotted in real time. At 0108 (UTC) we switched from using Camera 4 images in the real-time bug plotting to Camera 2; the focus identification routines on the Sun appear to work more efficiently on images from this camera and the volume imaged by Camera 2 is considerable greater than that of Camera 4.

The day dawned sunny and warm, with calm seas. By 1145 (UTC) we had completed 13 legs of the star, towyoing between the surface and 40 m. We conducted a deep towyo to near bottom following the turn onto Leg 14 and then resumed towyoing to 40 m. We started observing some kind of green or yellow material on the surface and in the upper water column; it was similar to pollen. Scott collected some of the material with a bucket. At 1434 (UTC) we completed Leg #16 and decided to look for an especially high concentration of Calanus in which to deploy a drifter, drogued to 30 m to tag the water parcel. We moved over to a location where we thought we had seen high concentrations of Calanus. On our way we passed through several pollen slicks. We selected 41° 21.808, 68°'a1 54. as our target location. We ended VPR 10 at 1625 (UTC) in order to deploy the drifter (Table 3). At 1645 (UTC) we had ship drills.

We deployed drifter #36 first but the drogue tether became tangled upon deployment and we had to retrieve it. One of the antennae broke against the side of the ship during retrieval and we had to switch to Drifter #21. We deployed drifter #21 at 1723 (UTC). The VPR was deployed for VPR 11 at 1818 (UTC) (Table 3). The purpose of VPR 11 was to document diel changes in the distribution of plankton in that water parcel. VPR 11 consisted of slow (4 kt) deep (surface to bottom) towyos around the drifter, with a radius of about 0.5 nm. We saw Calanus at the thermocline, coincident with the fluorescence and attenuation maxima, with a second peak at depth at the transition between the cold mid depth water to the warmer, saltier bottom water. We passed through a pollen slick at 2115 (UTC) and noted both temperature (13° to 16°) and salinity (32 to 31.5 psu) changes in the near surface water on the flow-through system on board ship as we traversed through the slick. Dolphins followed the VPR, the weather was sunny and calm, and a basking shark was spotted at 2125 (UTC). We saw two layers of both Calanus and pteropods: one at 30-40 m and the other at 80-90m. These distributions coincided with pycnoclines at both 30 and 90 m. We started to notice that the instrument was being offset to the side by strong currents at depth; the ADCP showed very low surface currents but a strong current at depth.

June 21, 1997 - Saturday

0000 (UTC) The fog had returned. We were still towyoing from the surface to 20 m above the bottom around the drifter. At 0100 (UTC) we noticed a strong smell of plankton, coincident with high abundances of both Calanus and pteropods near the surface (this distribution was seen in the real time plot of Calanus and pteropod distributions). The fog had lifted and the moon appeared to be nearly full. It was a glorious evening, with moonlight and hardly a ripple on the sea. Dolphins were observed sporadically, once riding the bow wave. At 0335 (UTC) we were in another patch of plankton. At 0352 (UTC) we terminated VPR 11 in order to conduct ring net tows to collect live pteropods for Scott Gallager. Approximately 4 ring net tows were conducted, yielding good hauls of plankton and plenty of pteropods. By 0913 (UTC), the fog had returned and it was necessary to furnish drifter positions to the bridge every 5 minutes.

Dawn arrived foggy. By 1229 (UTC) the fog had started to lift and the seas were fairly calm, although not as smooth as last evening. The Tans Vector was not functioning hence the ADCP is not getting Ash-gyro fixes, despite Jan’s best efforts. We suspected this was because we were sailing in 1 nm diameter circles. By 1422 (UTC) high currents were observed at depth again. The drifter has followed almost an entire tidal ellipse, traversing first to the south and now to the north. We thought we saw evidence of diel migration or redistribution of Calanus in the water column from a concentration at ~30 m at night to a distribution that extended throughout the water column during the day. We also saw a dramatic shift in water mass type throughout the water column in the salinity data which roughly corresponded to the night period which may obfuscate the diel signal. At 1815 (UTC) we ended VPR #12 and retrieved the drifter at 1838 (UTC). We headed for home with 420 hours of VPR towing completed, having achieved all of our objectives and more.

June 22, 1997 - Sunday

We arrived Woods Hole at 0530 local time. We unloaded all of our equipment and departed the dock by 0930 local time. R/V Endeavor was scheduled to depart mid-afternoon for Narragansett.

3. Preliminary Results

The cruise was extremely successful. All of the scientific objectives related to GLOBEC were met. Furthermore, we were able to document distributions of plankton in real time for the first time. In total, we towyoed the VPR for 212 hours at 8 knots covering a total distance of 1448 nautical miles. The locations of all VPR data collections are shown in Figure 1. We will be able to combine the plankton distributions with the ADCP velocities and the drifter tracks in indication of the flux of Calanus and Pseudocalanus through the region; preliminary considerations suggest that flux through the study region was primarily to the southwest and off of Georges Bank. There appeared to be little retention of water on the Bank even as shallow as 65 m. The data on plankton size and taxonomic composition as well as hydrography, light, fluorescence, and attenuation will provide further insights into the factors controlling plankton populations on Georges Bank.

The success of the cruise was attributable to several factors including good weather, reliable equipment, and the excellent support of the ship's officers and crew. We experienced only a few minor equipment problems which were easily rectified in a short period of time. The automatic identification system functioned properly and furnished us with the first real-time plankton distributions measured using the video plankton recorder.

3.1 North-South 24 Hour Flux Study (VPR 2,3)

A 24 hour study of the plankton distributions and ADCP vectors along a N-S transect located just to the east of the Schlitz mooring line was conducted to quantify flux of plankton and particles through this section over an entire tidal cycle. We completed 12 N-S surveys of the towyo transect in approximately 24 hours. The hydrographic data (Figs. 2-4) demonstrated warm, saline water (Slope Water) in the lower water column (> 50 m), a surface mixed layer, and cold pool water between along the southern half of the transect (bottom depth > 60 m) during the first 10 transects. Warm, higher salinity water was observed throughout much of the water column (>15m) at the southern end (bottom depth > 60 m) of the last two transects. In the northern portion (bottom depth <60m) had uniform temperature, salinity, and density. Fluorescence (Fig. 2, 5) showed variation in magnitude and depth. Generally, fluorescence was greatest at the southern end of the transects was concentrated either in the upper water column (<15m) or just below that depth interval (10-20m); fluorescence at the northern end of the transects was of lower magnitude and well-mixed throughout the water column. Attenuation (Fig. 2) showed a distinct pattern, being high in the upper 20 m across the entire length of the transects but low below 20 m at the southern end of the transects and high throughout the water column at the northern end. ADCP velocities (uncorrected for tides; Fig. 6) demonstrated the influence of the tidal ellipse.

High concentrations of Chaetoceros socialis colonies and Obelia medusae were observed along the transects, especially at the northern edge, in both deep and shallow water. At southern end of transects, Obelia, chain diatoms, and Chaetoceros were found in the mixed layer and down to the cold pool/warm water interface. Copepods were seen below the Chaetoceros layer and teardrop shaped marine snow at depth. At the northern end of the transect, there was a well mixed water column with Chaetoceros throughout and high concentrations of medusae and teardrop shaped marine snow near the bottom. Copepods were seen below the Chaetoceros layer and teardrop shaped marine snow at depth.

3.2 Diel Migration Study (VPR 4)

A 24 hour survey around three drifters was conducted to identify diel changes in the vertical distribution of plankton, especially copepods. Our initial impression was that there was no diel change in the distribution of copepods but that fluorescence varied exhibited diel variation, especially in the upper 10 m (note higher fluorescence during periods of darkness; Fig. 7). Variation in hydrographic characteristics was observed as the ship moved from north to south around the drifter array; water to the south (>50 m bottom depth) was warm and saline at depth and throughout the water column while water to the north was colder and fresher.

3.3. Survey of retention/loss region (VPR 5,6)

A grid covering the retention/loss region was sampled over a continuous 2+ day period (Figs. 1 and 8; Table 3). Three sets of drifters launched at 65 m, 75 m, and 85 m water depth were tracked continuously (Fig. 8). One of the drifters that had been launched at 85 m tracked due south and was retrieved early. Four additional drifters were retrieved after three days and tracked to the southwest parallel to the isobaths. The sixth drifter, launched at 65 m and drogued to 30 m, did not move as quickly to the southwest. This drifter remained deployed for 5 days, until it became clear that the drifter was moving to the southwest and would not divert to the north.

Hydrographically, the southern portion of the grid contained warm, high salinity water at depth, cold pool water at mid-depths, and mixed layer water at the surface (Figs. 9 and 10). The northern portion of the grid contained colder, fresher water of more uniform temperature and salinity characteristics. High fluorescence was observed in the upper water column at the southern end of the transect and at mid-depths at the western end at about 70 m bottom depth (Fig. 11). Uniform fluorescence was observed throughout the water column in the northeastern corner of the grid. Preliminary observations suggest that much of this fluorescence was because of Chaetoceros socialis colonies.

Acoustic Doppler current profiler data from the grid will be detided and used to calculate fluxes of plankton, especially Calanus, through the region.

3.4 East-West 24 Hour Flux Study (VPR 7)

A 24 hour study of the plankton distributions and ADCP vectors along a E-W transect located just to the north of the Schlitz E-W mooring line was conducted to quantify flux of plankton and particles through this section over an entire tidal cycle (Fig. 1). We completed 12 E-W surveys of the towyo transect in approximately 25 hours (Table 3).

Hydrographically, temperature and salinity did not show substantial variation over much of the water column. Surface layers were warmer and fresher on the western ends of the transect lines (Figs. 12 -13) Fluorescence was greater at the eastern ends of the transect lines on the edges of Georges Bank where the bottom depth was < 50 m (Figs. 12, 15). Elevated fluorescence also was observed at the western ends of lines 7-10 (Line 1 is located at the bottom of the page).

We were able to describe the abundance of C. socialis colonies in real time using automated image detection and identification routines. Elevated abundances of C. socialis were associated with very dense water at depth on western end of lines 5-6 and 11-12 (Fig. 16; color scale is as for Figure 14; ranges associated with colors shown in each plot title). Examination of acoustic Doppler current profiler velocities from the transects indicated that this dense water was associated with flow convergent flow from the southwest/northwest along those lines (Fig. 17).

3.5 Survey of Eastern Side of the Great South Channel and the edge of Georges Bank (VPR 8)

A short survey of the western edge of Georges Bank and the eastern GSC was conducted to identify water masses and potential flux of plankton from the southwestern corner of the Bank around to the north in this region (Fig. 1; Table 3). A distinct transition in water mass characteristics was observed moving from south to north across the surveyed region; the southern portion was characterized by a well-mixed water column while the northern portion as stratified both thermally and also in terms of salinity (Figs. 18, 19). Stratification increased moving from east to west off the Bank and into the Great South Channel. Fluorescence was fairly homogenous throughout the water column in the southern section but exhibited marked variation, with greatest values at approximately 10 m and coincident with the base of the thermo- halo-cline in the northern portion (Fig. 20). Chain diatoms and marine snow were abundant throughout the well-mixed water column to the south but absent in the stratified water along the western ends of transect lines to the north. Calanus was observed at depth in the stratified water of the GSC.

3.6 Chevron Survey of the Great South Channel (VPR 9)

A "chevron" shaped survey was conducted in the "SCOPEX" region of the GSC to identify locations of patches of Calanus that would be useful for a survey of patch dynamics (Fig. 1; Table 3). The GSC was characterized by a stratified water column, with a shallow pycnocline at approximately 20 m (Fig. 21, 22). Fluorescence was greatest at the pycnocline and varied markedly at different locations across the transects. Lowest fluorescence was seen on the shallow portions of the Bank to the east (Fig. 23). Calanus was observed in varying concentration across the transect lines, with greatest abundances seen at the western edge near the shipping lanes. Vertically, Calanus was found at the pycnocline in association with the fluorescence maximum.

3.7 Patch Mapping Study (VPR 10)

A patch mapping study was conducted at a location in the GSC where elevated abundances of Calanus had been observed during the chevron survey (Fig. 1; Table 3). The study region remained stratified thermally throughout the survey (Fig. 24), however marked variation in salinity was observed with a transition varying salinity throughout the water column to homogenous, and fresh salinity during the latter portion of the survey (Fig 25). A marked fluorescence maximum was observed at the base of the thermocline (Fig. 26).

3.8 Diel Vertical Migration Study in the GSC (VPR 11,12)

During the last 24 hours of the cruise, we conducted a study of the diel vertical distribution of plankton in a parcel of water marked with a drifter drogued to 30 m (Fig. 27). A region with elevated abundances of Calanus was located as the starting point of the study. Although variations in hydrographic properties were observed during the course of the survey (see salinity in Fig. 24), variations were much less than had been seen during VPR 10. The water column remained stratified with a marked thermocline at 20 m (Fig. 28). Fluorescence and attenuation maxima were observed coincident with the base of the thermocline. Calanus (Fig. 29) were observed both at depth (~80 m) and at the thermocline (note that the color scale in this figure is incorrect; relative abundances range from low (blue) to high (red)). During night (see light in Fig. 28), Calanus was observed throughout the upper 20 m. Note also that the vertical distribution of salinity, density, attenuation, and Calanus remained an approximately constant distance from the bottom and hence followed the shape of the bottom topography.

3.9 Pteropod collection

Pteropods were collected using a ring net part way through the diel vertical migration study for S. Gallager (Table 3).

4.0 Cruise Participants

Scientific Personnel

Carin Ashjian WHOI, Co-Principal Investigator Chief Scientist

Cabell Davis WHOI, Co-Principal Investigator Scientist

Scott Gallager WHOI, Co-Principal Investigator Scientist

Xiaoou Tang WHOI Postdoctoral Investigator

Craig Lewis Dartmouth College Postdoctoral Investigator

Andy Girard WHOI Laboratory Assistant

Betsy Broughton NMFS Guest Scientist

Erik Anderson St. Francis Xavier University Guest Student

Rob Granucci Iona College Guest Student

Sarah Foster Dalhousie University Guest Student

Jan Szelag URI Marine Technician

Ship's Officers and Crew

Thomas Tyler Captain

Everett McMunn Chief Mate

Steve Vetra Second Mate

Jack Buss Boatswain

Glenn Prouty Able-Seaman

Dick Foley Able-Seaman

Dave Rocha Able-Seaman

Richard Smith Chief Engineer

Jim Cobleigh Assistant Engineer

Tim Varney Assistant Engineer

Dan Butler Cook

Dave Philbrick Messman

5.0 Scientific, Crew Watch Schedule

0000-0400 Davis, Lewis, Foster, Vetra, Rocha

0400-0800 Gallager, Girard, Anderson, McMunn, Prouty

0800-1200 Ashjian, Broughton, Granucci, Tyler, Foley

Table 1. Waypoints for the Zig-Zag survey


Latitude (°N)

Longitude (°W)


40° 30.00'

68° 15.0'


40° 37.50'

68° 20.0'


40° 45.00'

68° 24.5'


40° 51.50'

68° 30.0'


40° 50.75'

68° 42.0'


40° 49.50'

68° 55.5'


40° 48.50'

69° 09.0'


40° 39.00'

69° 03.5'


40° 29.00'

68° 57.0'


40° 17.00'

68° 51.0'


40° 21.50'

68° 39.5'


40° 26.00'

68° 27.5'

Table 2. Waypoints for the Great South Channel "Chevron" Survey


Latitude (°N)

Longitude (°W)


41° 23'

68° 38'


41° 26.5'

68° 48'


41°'a1 28'



41° 23'

69° 16.5'


41° 17'

69° 14'


41° 25'

68° 57'


41° 20'

68° 50'


41° 17'

68° 43'


41° 12'

68° 47'


41° 16'



41° 12'

69° 10'


41° 08'

69° 06'


41° 10'



41° 08'

68° 53'

Table 3. GLOBEC Data Log

Appendix 1.

R/V Endeavor Cruise #302

Cruise Plan (Subject to Change)

1. Monday, June 9, ~14:30

Endeavor arrives Woods Hole. Loading of cruise gear, winch, and fish commences.

2. Tuesday, June 10, TBA. Endeavor departs Woods Hole. Test tow of VPR conducted in Vineyard Sound.

3. Proceed to Waypoint #3. Deploy VPR. Conduct 25 hour tow at ~8 knots back and forth between Waypoints #3 and #1 (see Table 1). Recover VPR.

4. Drifter Deployments: Deploy 2 drifters (10 m and 30 m drogues) at each of three locations at 65, 75, and 85 m. Positions will be located to the west of the NS Schlitz mooring line, along a line parallel to the mooring line.

5. Diel Study around drifters: Deploy VPR and tow for 24 hours at ~4 kts. around at least 2 of the drifter pairs.

6. Zig-Zag Survey: Recover VPR and steam to Waypoint #1. Launch VPR. Steam at 10 kts in a zig-zag survey according to the following sequence:

Waypt. #1 to #4

Waypt. #4 to #12

Waypt. #12 to #5

Waypt. #5 to #11

Waypt. #11 to #6

Waypt. #6 to #10

Waypt. #10 to #7

Waypt. #7 to #4

Waypt. #4 to #8

Waypt. #8 to #3

Waypt. #3 to #9

Waypt. #9 to #2

Waypt. #2 to #10

Waypt. #10 to #1

NOTE: We may need to break off from the survey temporarily to recover drifters if they go south off the shelf or to maintain drifters (batteries).

7. Recover drifters that are leaving the sampling region

8. Conduct 25 hour tow at ~8 knots back and forth between Way point #6 and 40° 51.25', 68° 34'.

9. Recover Drifter #234.

10. Great South Channel Zig-Zag: Zig Zag survey of the western side of the Great South Channel in the following sequence:

40° 51.25 68° 34

41° 00 69° 00

41° 06 68° 36

41° 20.75 69°

41° 21.25 68° 28

41° 35.25 68° 55.5

41° 34 68° 24.5

(Note: last position located to the north of the region closed to fishing)

11. Return to the closed area and follow a chevron pattern E-W across the Great South Channel, between the Bank to the east and the shipping lanes to the west.

12. Return to an as yet unidentified position (Calanus mother lode) in the GSC and conduct a patch study by towyoing in a star pattern.

13. Launch drifter at position of high Calanus abundance. Conduct diel study by towyoing VPR around the drifter for a complete diel cycle.

14. Collect pteropods in the GSC at a location determined by Scott Gallager (ring net sample).